An object of mass 4 kg is moved on a horizontal surface with a constant acceleration of 0.2m/s^2 , by a horizontal force of 6 Newtons. Find the friction force acting on the object and the coefficient of friction.



1
Expert's answer
2022-03-21T12:41:55-0400

Net force

"F=ma"

Friction force

"F_1-F=F_f"

Coefficient of friction

"f=\frac{F_f}{R}"

"R=mg"

Calculating

"F_f=6-4\cdot 0.2=5.2"

"f=\frac{5.2}{4\cdot 10}=0.13"

Answer Friction force 5.2 newton, coefficient of friction 0.13